I need help with 2/3 three problems:
1) Find the probability of the result using the normal curve approximation to the binomial distribution. On a hospital floor, 60 patients have a disease with a mortality rate of 0.1. Five of them die.
2) Find the probability fo the result using the normal curve approximation to the binomial distribution. A battery company has found that the defective rate of its batteries is 0.03. Each day, 400 batteries are randomly tested. On Tuesday, 14 are found to be defective.

Answer by stanbon(48568)

1) Find the probability of the result using the normal curve approximation to the binomial distribution. On a hospital floor, 60 patients have a disease with a mortality rate of 0.1. Five of them die.
------
The mean is np = (60)(0.1)
The standard dev. = sqrt(npq)
Using a TI caluculator you get normalcdf(4.5,5.5,60*0.1,sqrt(60*0.1*0.9))=0.1555..
----------------
2) Find the probability fo the result using the normal curve approximation to the binomial distribution. A battery company has found that the defective rate of its batteries is 0.03. Each day, 400 batteries are randomly tested. On Tuesday, 14 are found to be defective.
----------
Mean = 400*0.3
Std = sqrt(400*0.03*0.97)
normalcdf(13.5,14.5,400*0.03,sqrt(400*0.03*0.97)) = 0.09824...
----------
Cheers,
Stan H.